OpenStudy (anonymous):

cos^[(pi/2)-x]/ cos x Match the trig expression with one of the following a) csc x b) tan x c) sin^2 x d) sin x tan x e) sec^2 x f)sec^2 x + tan^2 x

OpenStudy (anonymous):

cos^[(pi/2)-x]/ cos x = sin x/ cos x = tan x Hence option b is correct

OpenStudy (anonymous):

cos^2[(pi/2)-x]/cos x there was a typo is the answer still the same?

OpenStudy (jdoe0001):

well, the answer will still hold, with just an extra sine function so $$ \cfrac{cos^2\pmatrix{\frac{\pi}{2}-x}}{cos(x)}\\ \implies \cfrac{sin^2(x)}{cos(x)} \implies sin(x)\cfrac{sin(x)}{cos(x)} \implies sin(x)tan(x) $$

OpenStudy (anonymous):

@jdoe0001 could you explain how you got the first part?